(a) With what speed must a ball be thrown vertically from ground level to rise to a maximum height of 36 m? (b) How long will it be in the air? (a) Number[26.56 (b) Number24.11 UnitsT m/s UnitsT s
Answer #1

a)v^2=u^2+2as

v=0

a=-10m/s^2(since it is due to gravity in downward direction)

s=36m

u=26.56m/s

b)v=u+at

0=26.56-10t

Substituting values from above.

t=2.656s